OpenStudy (anonymous):

A circle has a diameter of ten millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth. A.fifteen point seven millimeters; seventy-eight point five square millimeters B.thirty-one point four millimeters; seventy-eight point five square millimeters C.fifteen point seven millimeters; three hundred fourteen point zero square millimeters D.thirty-one point four millimeters; three hundred fourteen point zero square millimeters

jhonyy9 (jhonyy9):

diameter = 10 mm so this mean that radius = diameter /2 = 5 mm circumference = 2*pi * r = ... area = pi*r^2 = ... hope this will help you